# Formalising the intentional stance: attributing goals and beliefs to stochastic processes

### v1

@@ 7 Limitations and future work
In terms of limitations, one crucial missing component of this work is Dennett’s idea that for some systems — the co-called intentional systems — the intentional stance is much more productive than for others, and in particular more productive than the physical stance. Our current framework only says which teleo-environments can be consistently attributed as beliefs; it doesn’t currently offer any way to say which belief attributions are ‘better,’ or more natural or pragmatic for a given agent’s behaviour. After all, in Theorem 6.3.2 we showed that one can always attribute a set of beliefs that say, in effect, that the agent believes it must act in exactly the way it does act, otherwise it will fail its goal; but we suspect it will rarely if ever be fruitful to explain a real agent’s behaviour in terms of such a belief. We suspect the answer might lie in bounded rationality, which we have only made a start on considering here.
@@ ACKNOWLEDGEMENT
This article was produced with financial and technical support from Principles of Intelligent Behaviour in Biological and Social Systems (PIBBSS), and Simon McGregor’s work was supported with a scholarship grant from the Alignment of Complex Systems Research Group (ACS) at Charles University in Prague. Nathaniel Virgo’s work on this publication was made possible through the support of Grant 62229 from the John Templeton Foundation. The opinions expressed in this publication are those of the author(s) and do not necessarily reflect the views of the John Templeton Foundation.
The authors would also like to thank Martin Biehl for feedback on the manuscript and Robert Obryk for fruitful discussions about technical details.

### v2

@@ 5 Limitations and future work
In terms of limitations, one crucial missing component of this work is Dennett’s idea that for some systems — the co-called intentional systems — the intentional stance is much more productive than for others, and in particular more productive than the physical stance. Our current framework only says which teleo-environments can be consistently attributed as beliefs; it doesn’t currently offer any way to say which belief attributions are ‘better,’ or more natural or pragmatic for a given agent’s behaviour. After all, in Theorem Theorem we showed that one can always attribute a set of beliefs that say, in effect, that the agent believes it must act in exactly the way it does act, otherwise it will fail its goal; but we suspect it will rarely if ever be fruitful to explain a real agent’s behaviour in terms of such a belief. We suspect the answer might lie in bounded rationality, which we have only made a start on considering here.
@@ ACKNOWLEDGEMENT
This article was produced with financial and technical support from Principles of Intelligent Behaviour in Biological and Social Systems (PIBBSS), and Simon McGregor’s work was supported with a scholarship grant from the Alignment of Complex Systems Research Group (ACS) at Charles University in Prague. Nathaniel Virgo’s work on this publication was made possible through the support of Grant 62229 from the John Templeton Foundation. The opinions expressed in this publication are those of the author(s) and do not necessarily reflect the views of the John Templeton Foundation. The authors would also like to thank Martin Biehl for feedback on the manuscript.